OpenStudy (anonymous):

What are the roots of the following polynomial equation? (x + 5)(x + 2i)(x – 5i) = 0

OpenStudy (across):

They're staring you right in the face!

OpenStudy (anonymous):

Set each piece equal to zero, via the zero product principle. Then, according to the complex conjugates theorem, you know that if -2i is a solution then...

OpenStudy (anonymous):

i got -2i , and 5i

OpenStudy (across):

There's one more.

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

actually... there are a total of 5 zeros

OpenStudy (across):

Do enlighten us!

OpenStudy (anonymous):

see my earlier post... complex conjugates theorem tells us that if -2i is a root of a polynomial equation, then +2i is ALSO a root.

OpenStudy (ledah):

Example 1. Let P(x) = 5x³ − 4x² + 7x − 8. Then a root of that polynomial is 1, because P(1) = 5· 1³ − 4· 1² + 7· 1 − 8 =5 − 4 + 7 − 8 =0 It is traditional to speak of a root of a polynomial. Of a function in general, we speak of a zero. This is just an example I swiped off a website but maybe it will help. Hope it does.

OpenStudy (across):

\((x + 5)(x + 2i)(x – 5i) = 0\) is a degree-3 polynomial. The fundamental theorem of algebra, basically, states that a polynomial of degree \(n\) has \(n\) roots. Which one is it?

OpenStudy (anonymous):

It's not... it's a 5th degree polynomial with 1 real root and 4 imaginary roots. This means it has only one x-intercept.

OpenStudy (across):

How it it a degree-5 polynomial?

OpenStudy (anonymous):

there are 3

OpenStudy (anonymous):

if you expand there is no way to get a higher than x^3 term.

OpenStudy (anonymous):

Look up "Complex Conjugates Theorem." It states that if 3 + i is a root of an equation, so is 3 - i. So, if you're only solving the equation listed, you're correct (there are only 3 terms listed in the product so it seems that there are only 3 solutions). However, as the question asked what are the roots of the equation, you must also consider the conjugates of any complex root.

OpenStudy (anonymous):

imaginary roots always come in pairs and i think 2i is the pair of 5i

OpenStudy (anonymous):

i believe the answer is more simple then where over looking it, -5,-2i,5i i think we just have to wright the inverse of the in the ( )

OpenStudy (across):

You got confused.

OpenStudy (anonymous):

No, that's not right, timo86m. 2i must pair with -2i and 5i must pair with -5i.

OpenStudy (across):

The Complex Conjugate Theorem requires that the polynomial has real coefficients. If you expand this thing out, you'll have complex ones.

OpenStudy (ledah):

This is an elaborate debate.

OpenStudy (across):

Guys, this is why I always recommend you take everything you get told in here with a huge grain of salt.

OpenStudy (anonymous):

I am with across on this one

OpenStudy (anonymous):

here is what you get when you expand it x^3-(3*I)*x^2+10*x+5*x^2-(15*I)*x+50

OpenStudy (anonymous):

Usually when a question is asked about roots in a HS class where some are imaginary, they are testing an understanding of Complex Conjugates theory... And this is coming from a veteran math teacher of 10 years. I guess it would be helpful to know if Complex Conjugates were discussed in this chapter or unit.

OpenStudy (anonymous):

ok even maple tells me {x = -5}, {x = -2*I}, {x = 5*I}

OpenStudy (across):

Clearly, that is not being tested here today.

OpenStudy (anonymous):

I didn't think it was clear and I'm sorry if I confused the discussion in any way. I was trying to help and nothing I've explained is wrong. If this is an Algebra 2 class, then I'm probably over-explaining. If this is precalc, then this would be part of the discussion... at least it is in the school where I teach.

OpenStudy (anonymous):

it doesn't matter the class the point is it has 3 zeros. Do you finally get that?

OpenStudy (across):

I know where you're coming from, and I totally agree with your bringing of the Complex Conjugate Theorem. I just don't want @Tea-Is-God thinking that a degree-3 polynomial has 5 roots. :)

OpenStudy (anonymous):

i think by his logic (x+2*I)*(x-5*I) would have 4 roots.

OpenStudy (anonymous):

thank you for the spirited debate, @across. @timo86m, there is more to my point than you're realizing. your comment and tone with "do you finally get that" was unnecessary given that i was only trying to help.

OpenStudy (across):

You're welcome. And @timo86m, that wasn't necessary at all.

OpenStudy (anonymous):

No i mean it in a way do you get it? it is literal question. Sorry it sounded bad.